OpenStudy (paradise):

Help me guys!! Solve by the elimination method. 2x – 5y = 1 3x – 4y = -2

OpenStudy (anonymous):

first you would isolate x in the first equation

OpenStudy (anonymous):

1st add a*(eq1) to eq 2 in such a way that it makes the 3x component 0 x

OpenStudy (anonymous):

3x+ a*2x=0 a*2x=-3x a=-3x/2x xs cancels so a=-3/2 or if you can remember a is always -( x coefficient of eq1/ x coefficient of eq2 )

OpenStudy (anonymous):

2x – 5y = 1 3x – 4y = -2 \/ -3/2 r1+r2 2x- 5y=1 0x-2/3y=-4/3 \/ now we must make y into 1 a*-2/3=1 -> the reciprocal -3/2 so multply r2 by it 2x- 5y=1 y =2 we now know y = 2 so we can back substitude into row 1 2x- 5(1)=1 2x-5=1 2x=6 x=3 so y=2 x=3

OpenStudy (paradise):

I got x=-2 and y=-1

OpenStudy (anonymous):

oh then idk lol let me check on a calculator

OpenStudy (anonymous):

you are right hmm. maybe i did an arithmetic error sorry.

OpenStudy (anonymous):

steps are right tho ax+by=c dx+ey=f get dx to equal 0 by adding some multiple of r1 to r2 ax+by=c ab+ey=ac+f get ab+ey to equal 1 by adding a multiple of r2 y=ac+f+a*somemultiple y= ac+f+a*somemultiple = n where n is some number you plug in n into first row and solve for x.